# Balancing Caregiving and Self-Care: Exploring Mental Health Needs of Alzheimer’s and Dementia Caregivers

### v1

@@ 1. Introduction
The mental health burden borne by family caregivers underscores the critical need for effective support systems and interventions to help them cope with the psychological demands of caregiving (Coon and Evans, 2009; Harper, [n. d.]). However, the specific mental wellbeing needs of AD/ADRD caregivers, and how these needs evolve throughout the caregiving journey, remain less understood. From a human-computer interaction (HCI) and computer-supported cooperative work (CSCW) perspective, there is also a gap in understanding the specific technology and collaboration needs of caregivers. In particular, caregiving is inherently multi-layered and collaborative, involving direct interactions between caregivers and recipients while also requiring coordination with family members, healthcare providers, government agencies, community resources, and supportive technological tools (Schulz et al., 2016; Robinson et al., 2021).
Prior work in the HCI/CSCW space has explored technologies such as smartphone applications and virtual assistants to facilitate collaboration between caregivers and care recipients, supporting tasks like medication management and routine coordination (Kim et al., 2024; Meyerhoff et al., 2022; Shreve et al., 2016; Boessen et al., 2017; Kourtis et al., 2019). However, the role of technology in supporting caregivers’ mental health remains largely underexplored.
Together, given the global increase in AD/ADRD, the growing reliance on informal and family caregivers (Nichols et al., 2022; Better, 2023), and the promises of HCI and personal health informatics in supporting mental health (Slovak and Munson, 2024; Epstein et al., 2020), it is essential to develop a deeper understanding of the caregiving journey and how technology can better support family caregivers’ mental wellbeing. As people’s needs evolve, so do their expectations of technology. Recent work in personal informatics has highlighted the need to design tools that not only understand individuals’ goals but also adapt as those goals evolve, providing better support (Munson et al., 2020; Schroeder et al., 2019; Sefidgar et al., 2024). As an AD/ADRD caregiver’s needs evolve over time, from seeking information about the condition to learning how to care to discovering effective lifestyle adaptations, their expectations from technology also evolve. Adopting a personal health informatics perspective for technology design, the key open question is how we can design mental wellbeing technologies that offer tailored and timely support to meet the unique needs of family caregivers? Accordingly, our research investigates the evolving mental wellbeing needs and concerns of AD/ADRD caregivers [^1] Since our work focuses on family caregivers, the term ’caregivers’ will refer to family caregivers unless otherwise noted.. Our work is guided by the following research questions (RQs):
@@ 3. RQ3:
We conducted 25 semi-structured interviews with family caregivers of AD/ADRD individuals in the U.S. The interviews focused on caregivers’ daily routines, mental health challenges, coping strategies, and the technologies they use. We analyzed the interview data using inductive qualitative coding, followed by thematic analysis (Braun and Clarke, 2019) to identify key themes catering to our RQs. Our findings indicate that caregivers experience multifaceted mental health concerns; we categorize these concerns into cause and effect and identify three stages of evolution for caregivers’ mental health needs. Participants expressed that while they realize the need to socialize and seek professional support to manage their mental wellbeing, they often feel constrained by both time and space to simultaneously balance the demands of professional, personal, and caregiving responsibilities. They reflected on the self-care practices they adopt and the external support they seek to manage caregiving and wellbeing concerns. In this regard, participants also shared the role played by a variety of technologies, ranging from monitoring and medication management to social belonging and support. We found that caregivers hold both optimism and skepticism about technologies, primarily around computing technologies such as online platforms, smart devices, and AI chatbots. For instance, while some participants expressed optimism and excitement about the timely and personalized responses offered by AI, others remained skeptical regarding the accuracy of the information and the absence of human interaction. Additionally, we found the key barriers caregivers face when using technology, and underscore the need for more affordable, accessible, and personalized mental health support tools.
@@ 2.1. Alzheimer’s Disease and Related Dementias (AD/ADRD): Condition and Caregiving
Alzheimer’s Disease and Related Dementias (AD/ADRD) is a family of neurodegenerative conditions that progressively worsens with no definitive cure (Kim et al., 2021). It remains a major public health concern, ranking as the fifth-leading cause of death among Americans aged 65 and older (Better, 2023). Projections indicate that by 2050, approximately 14M individuals in the U.S. and 152M worldwide will be living with AD/ADRD (Nichols et al., 2022). The caregiving for AD/ADRD is predominantly undertaken by family members and informal caregivers within the home setting (García-Alberca et al., 2011; Grunfeld et al., 1997). In 2022 alone, the unpaid caregiving provided by family members was valued at approximately \$339.5B USD (Better, 2023).
Additionally, family caregivers frequently experience financial strain (Gibson et al., 2014; Griffiths and Bunrayong, 2016; Unbehaun et al., 2018) and difficulties in planning for future crises (Millenaar et al., 2018; Vaingankar et al., 2013; Wawrziczny et al., 2017). Many also struggle with maintaining their own physical and mental wellbeing (Coon and Evans, 2009; Hughes et al., 2014; Wawrziczny et al., 2017). Therefore, this caregiving effort comes with significant personal costs, including a heightened risk of emotional distress and adverse mental and physical health outcomes for caregivers (Harper, [n. d.]; García-Alberca et al., 2011; Grunfeld et al., 1997; Better, 2023). In particular, caregivers of AD/ADRD individuals often face higher levels of stress compared to those caring for individuals with other conditions (Association et al., 2005). Without sufficient training or support, they encounter numerous challenges, such as managing the evolving symptoms of the care recipient (Rettinger et al., 2020; Ruitenburg et al., 2024), providing supervision (Queluz et al., 2020), and making complex medical decisions regarding comorbid conditions (Lee et al., 2022).
@@ 3.1. Participants and Recruitment
We recruited our participants primarily through social media. We first contacted the moderators of different online communities catering to AD/ADRD-related discussions on Reddit (r/alzheimers, r/dementia, r/dementiaresearch, r/ParentsWithAlzheimers, etc.) and AlzConnected (alzconnected.org), by briefly describing our research and asking if they were okay with recruiting from their respective platforms. Then, in online communities—where we received moderators’ approval—we posted our recruitment flyer with an interest form that included a demographic survey questionnaire (age, sex, race, U.S. state) and their role as a caregiver. This interest form helped us target and screen participants who are 1) 18 years or older, 2) current/former caregivers for AD/ADRD, and 3) residing in the U.S. We received 293 responses to our interest form over a period of two months between August and October 2024, and we invited a subset of participants to maximize diversity and balance across answers to the caregiving role and tenure. This led to a final set of 25 participants who consented to participate in the study, and we interviewed them. We note that although our recruitment flyers were posted on social media, three of our participants did not actively participate on these platforms, and were rather referred to by others to express their interest in participating in our study. Each participant was compensated with Amazon gift vouchers of \$25 USD.
Table 1 summarizes the demographic and caregiving role and tenure of the participants. Of the participants we interviewed, 76% (19 out of 25) are current caregivers, and we note a diversity of participants across age group, number of years in caregiving, race, education, and occupation. Before the interviews, participants were provided with the Rapid Caregiver Well-being Scale (R-CWBS) (Tebb et al., 2013) along with the consent form. R-CWBS is a validated short-form rapid assessment instrument to infer key areas of support a caregiver needs (Tebb et al., 2013). Here, each question is rated on a Likert-scale between 1 (Rarely) and 5 (Usually), and lower scores indicate a need for greater support. Table 2 provides a summary of participants’ responses to this survey, showing that although our pool of participants was mostly regular with taking care of personal daily activities, the other questions received a variety of responses.
Table 1. Summary of participants, including Type (Current/Former Caregiver), Years of Caregiving (Ys.), Care Recipient, Age, Gender, Race, Education, and Occupation. Some caregivers started their caregiving for a family member, and later transitioned into professional caregiving. Serial Caregivers are marked with an ’*’ next to their ID.
@@ 3.2. Interview Procedure
In particular, we shared our screen with participants and asked them to rate various caregiving-related concerns—1) disruptive patient behaviors (Swearer et al., 1988; Kunik et al., 2003; Desai et al., 2012), 2) insufficient support systems (Schulz and Martire, 2004; Jennings et al., 2015; Park et al., 2018), 3) doubt in self-efficacy (Tang and Chan, 2016; Gallagher et al., 2011; Crellin et al., 2014), 4) emotional wellness issue (Aminzadeh et al., 2007; Caputo, 2021; Gallego-Alberto et al., 2022), 5) relationship management (Ray, 2022; Vroman, 2019), 6) compassion fatigue (Day and Anderson, 2011; Day et al., 2014; Pérez et al., 2022), 7) no time for self-care (Waligora et al., 2019; Oliveira et al., 2019; Wang et al., 2019), and 8) burnout (Alves et al., 2019; Truzzi et al., 2012; Takai et al., 2009). These eight concerns were obtained from prior literature as significant challenges for AD/ADRD caregivers. This approach enabled us to situate our findings with existing knowledge while also allowing for the discovery of novel insights specific to our participants’ experiences. We guided participants through each concern, providing literature-based definitions, and asked them to rate their level of concern on a scale from 1 (not at all concerning) to 5 (very concerning). Additionally, we asked the participants to explain their ratings and think aloud about specific experiences related to these concerns. This approach helped build common ground and helped us gather deeper insights into participants’ mental wellbeing concerns. Table 3 summarizes participants’ responses to these prompts, where we find a high concern for most of the prompts. In our ensuing qualitative analyses (Braun and Clarke, 2019), we incorporated the deeper insights that they shared. Finally, we asked the participants about the use of technology in caregiving as well as in managing mental wellbeing. Participants were encouraged to share their thoughts on the technologies they used, what features they found useful, and any suggestions or concerns they had for improving their experience. This helped us understand participants’ concerns and desires about these technologies.
@@ 3.3. Data Analysis
All co-authors participated in reviewing the transcripts and engaged in an iterative process of open coding, where codes were grouped, initial subthemes were identified, and these subthemes were refined into higher-level themes. To further elaborate our process, all five co-authors were involved in the open coding process. The co-first authors led the coding effort, completing the majority of coding on raw interview transcripts, while the senior co-authors (who have extensive experience in qualitative research) coded two raw transcripts each during two two-hours hybrid (in-person and screen-sharing-based) co-working sessions. To ensure coherence in the analysis, we carefully reviewed and refined the themes. This process involved merging certain themes into broader categories, separating other themes into distinct categories, and discarding themes that were not directly relevant to our core research questions. Ultimately, we determined the final themes corresponding to each research question. While we allowed flexibility in our coding and theme development, we employed thematic analysis (Braun and Clarke, 2019) as our primary methodological approach, guided by prior research on social support in online dementia communities (Johnson et al., 2022). A random subset of 10 transcripts was selected for an initial round of open coding. We developed 400 codes in the first iteration, which were then discussed by the entire team and distilled into 851 codes. The new set of codes were applied and iterated on the remaining transcripts. We discovered an additional 451 codes from the remaining transcripts. The codes were then grouped into 12 higher-level and 63 lower-level themes, which aligned with our three research questions.
@@ 4.1.1. Intersecting Sources of Caregivers’ Distress
Time Constraints and Limited Personal/Self-care Time.  Participants expressed the continual struggle with time management due to caregiving demands, which often leaves little time for self-care. Many felt overwhelmed by having to prioritize caregiving tasks, often at the expense of their own wellbeing. P4 described caregiving as a highly time-consuming job, explaining that factors such as commuting, unexpected incidents, and the care recipient’s declining verbal abilities made it increasingly difficult to engage in other activities. P7 and P15 emphasized hat caregiving often requires attention 24/7, leaving no time for breaks. Some, like P19, even changed jobs to better accommodate caregiving responsibilities:
“As a school Superintendent, that was really stressful. I was responsible for taking care of thousands of people, and I couldn’t do that full-time while also taking care of my mom. So I decided to resign and take a job with the university where I can work from home.” —P19
@@
“I’ve stopped working, I lost my friends, like, I feel like I’ve sacrificed a lot for her.” —P3
@@ 4.1.2. Mental Health Impacts on Caregivers
The reality is that this is the only disease without a cure, and this disease is fatal, so we’re dealing a lot with anticipatory grief, which is what we talk about a lot. [..] And many times, people are not going to support groups until they’re like, literally in tears or at wits’ end. I see anger and frustration from men, and I see tears and physical demise from women. —P4
@@
“I feel increasing [caregiving] demands lead to physical and emotional burden, affecting my mental health. I became more concerned about burnout and maintaining my wellbeing.”—P21
@@
Overall, the caregiving-related concerns identified in prior research served as useful starting points for our conversations with participants (Table 3). These prompts helped guide the participants in recalling and narrating personal experiences that aligned with known themes, such as emotional distress, burnout, and lack of support. These concerns were strongly echoed in the participant narratives. Many participants described feeling overwhelmed by the constant demands of caregiving, often at the expense of their own wellbeing. In addition, the concerns surfaced in our study ultimately represented an assimilation of themes that emerged organically across diverse participant narratives. For example, discussions around financial burden and its impact on mental health, as well as the erosion of personal and social life while caregiving for individuals with AD/ADRD, arose naturally through these conversations. Additionally, issues such as relationship strain, guilt surrounding self-care, and limited family support emerged as significant concerns—ones that were not always directly framed as mental health issues in prior work. Notably, participants also reflected on the self-reflective positive aspects of caregiving—such as personal growth and emotional resilience—that extended beyond the more problem-oriented concerns highlighted in prior work.
@@ 4.2. The Evolution of Mental Health of Caregivers
“When I became a caregiver, I didn’t realize how much time and energy would be consumed by caregiving. I was in denial about how much this would impact me emotionally.”—P21
@@
“I started to like experience, feelings of burnout and stress. I found that I needed more emotional support, not just from my colleagues, but also from therapy or peer groups.”—P17
@@
“Now, I have no one to open up to. I’m not comfortable sharing my problems with someone I know. I feel they may judge me—or treat my mom differently when they visit our home.”—P3
@@
Late Stage: Emotional Exhaustion and Burnout.  Over the significant progression of time and the decline of care recipient’s condition, caregivers adapt to their caregiving duties almost like a full-time job. However, by this point, they are also overwhelmed with providing 24/7 care. This leaves them with no personal time to address their own needs. This late stage represents the peak of both physical and emotional exhaustion, with burnout becoming a frequent concern. Participants also described how they can only get little or no sleep, as P11 described they would almost “sleep with one eye open.” P10 described the demands of managing every aspect of the care recipient’s life, including handling bills and estate matters, coordinating constant medical appointments, and providing hands-on care and support.
@@
Experience/Journey of Serial Caregivers.  A key consideration to note about this three-stage process is that it primarily represents the typical experience of a first-time caregiver. For people who are serial caregivers (i.e., they have cared for multiple family members with AD/ADRD), such as P2, P16, P17, P21, and P22 in our study, the experience changes over iterations. That said, the rate at which technology is evolving also affects these experiences. For instance, P2 mentioned that having smart home technologies earlier, would have had a significant impact when they were caregiving two decades ago.

### v2

@@ 1. Introduction
The mental health burden borne by family caregivers underscores the critical need for effective support systems and interventions to help them cope with the psychological demands of caregiving (Coon and Evans, 2009; Harper, 2022). However, the specific mental wellbeing needs of AD/ADRD caregivers, and how these needs evolve throughout the caregiving journey, remain less understood. From a human-computer interaction (HCI) and computer-supported cooperative work (CSCW) perspective, there is also a gap in understanding the specific technology and collaboration needs of caregivers. In particular, caregiving is inherently multi-layered and collaborative, involving direct interactions between caregivers and recipients while also requiring coordination with family members, healthcare providers, government agencies, community resources, and supportive technological tools (Schulz et al., 2016; Robinson et al., 2021). Prior work in the HCI/CSCW space has explored technologies such as smartphone applications and virtual assistants to facilitate collaboration between caregivers and care recipients, supporting tasks like medication management and routine coordination (Kim et al., 2024; Meyerhoff et al., 2022; Shreve et al., 2016; Boessen et al., 2017; Kourtis et al., 2019; Theisz et al., 2025). However, the role of technology in supporting caregivers’ mental health remains largely underexplored.
Together, given the global increase in AD/ADRD, the growing reliance on informal and family caregivers (Nichols et al., 2022; Alz, 2024), and the promises of HCI and personal health informatics in supporting mental health (Slovak and Munson, 2024; Epstein et al., 2020), it is essential to develop a deeper understanding of the caregiving journey and how technology can better support family caregivers’ mental wellbeing. As people’s needs evolve, so do their expectations of technology. Recent work in personal informatics has highlighted the need to design tools that not only understand individuals’ goals but also adapt as those goals evolve, providing better support (Munson et al., 2020; Schroeder et al., 2019; Sefidgar et al., 2024). As an AD/ADRD caregiver’s needs evolve over time, from seeking information about the condition to learning how to care to discovering effective lifestyle adaptations, their expectations from technology also evolve. Adopting a personal health informatics perspective for technology design, the key open question is how we can design mental wellbeing technologies that offer tailored and timely support to meet the unique needs of family caregivers?
Accordingly, our research investigates the evolving mental wellbeing needs and concerns of AD/ADRD caregivers [^1] Since our work focuses on family caregivers, the term ’caregivers’ will refer to family caregivers unless otherwise noted.. Our work is guided by the following research questions (RQs):
@@ 3. RQ3:
We conducted 25 semi-structured interviews with family caregivers of AD/ADRD individuals in the U.S., focusing on their routines, mental health challenges, coping strategies, and technology use. Using inductive qualitative coding and thematic analysis (Braun and Clarke, 2019), we identified key themes aligned with our RQs. Participants reported multifaceted mental health concerns, which we categorized into causes and effects across three stages of evolving needs. While they recognized the importance of socializing and seeking professional support, many felt constrained by time and competing responsibilities. In this regard, they described both self-care practices and external support mechanisms they rely on. They also discussed the role of technologies—from monitoring tools to social platforms and AI chatbots—in supporting their wellbeing. Caregivers expressed mixed views: some were optimistic about the personalization and timeliness of AI support, while others were skeptical of its accuracy and lack of human connection. Finally, we found the key barriers to technology use, including cost, accessibility, and lack of personalization in current mental health tools.
@@ 2.1. Alzheimer’s Disease and Related Dementias (AD/ADRD): Condition and Caregiving
Alzheimer’s Disease and Related Dementias (AD/ADRD) is a family of neurodegenerative conditions that progressively worsens with no definitive cure (Kim et al., 2021). It remains a major public health concern, ranking as the fifth-leading cause of death among Americans aged 65 and older (Alz, 2024). Projections indicate that by 2050, approximately 14M individuals in the U.S. and 152M worldwide will be living with AD/ADRD (Nichols et al., 2022). The caregiving for AD/ADRD is predominantly undertaken by family members and informal caregivers within the home setting (García-Alberca et al., 2011; Grunfeld et al., 1997). In 2022 alone, the unpaid caregiving provided by family members was valued at approximately \$339.5B USD (Alz, 2024).
Additionally, family caregivers frequently experience financial strain (Gibson et al., 2014; Griffiths and Bunrayong, 2016; Unbehaun et al., 2018) and difficulties in planning for future crises (Millenaar et al., 2018; Vaingankar et al., 2013; Wawrziczny et al., 2017). Many also struggle with maintaining their own physical and mental wellbeing (Coon and Evans, 2009; Hughes et al., 2014; Wawrziczny et al., 2017). Therefore, this caregiving effort comes with significant personal costs, including a heightened risk of emotional distress and adverse mental and physical health outcomes for caregivers (Harper, 2022; García-Alberca et al., 2011; Grunfeld et al., 1997; Alz, 2024). In particular, caregivers of AD/ADRD individuals often face higher levels of stress compared to those caring for individuals with other conditions (Association et al., 2005). Without sufficient training or support, they encounter numerous challenges, such as managing the evolving symptoms of the care recipient (Rettinger et al., 2020; Ruitenburg et al., 2024), providing supervision (Queluz et al., 2020), and making complex medical decisions regarding comorbid conditions (Lee et al., 2022).
@@ 3.1. Participants and Recruitment
We recruited our participants primarily through social media. We first contacted the moderators of different online communities catering to AD/ADRD-related discussions on Reddit (r/alzheimers, r/dementia, r/dementiaresearch, r/ParentsWithAlzheimers, etc.) and AlzConnected (alzconnected.org), by briefly describing our research and if they were okay with recruiting from their respective platforms. Then, in online communities, where we received moderators’ approval, we posted our recruitment flyer with an interest form that included a demographic survey questionnaire (age, sex, race, U.S. state) and their role as a caregiver. This interest form helped us target and screen participants who are 1) 18 years or older, 2) current/former caregivers for AD/ADRD, and 3) residing in the U.S. We received 293 responses to our interest form over a period of two months between August and October 2024, and we invited a subset of participants to maximize diversity and balance across answers to the caregiving role and tenure. This led to a final set of 25 participants who consented to participate in the study, and we interviewed them. We note that although our recruitment flyers were posted on social media, three of our participants did not actively participate on these platforms, and were rather referred to by others to express their interest in participating in our study. Each participant was compensated with Amazon gift vouchers of \$25 USD.
Table 1 summarizes the demographic and caregiving role and tenure of participants. Of the participants we interviewed, 76% (19 out of 25) are current caregivers, and we note a diversity of participants across age group, number of years in caregiving, race, education, and occupation. Before the interviews, participants were provided with the Rapid Caregiver Well-being Scale (R-CWBS) (Tebb et al., 2013) along with the consent form. R-CWBS is a validated short-form rapid assessment instrument to infer key areas of support a caregiver needs (Tebb et al., 2013). Here, each question is rated on a Likert-scale between 1 (Rarely) and 5 (Usually), and lower scores indicate a need for greater support. Table 2 provides a summary of participants’ responses to this survey, showing that although our pool of participants was mostly regular with taking care of personal daily activities, the other questions received a variety of responses.
Table 1. Demographic information regarding the participants, including Type (Current/Former Caregiver), Years of Caregiving (Ys.), Care Recipient, Age, Gender, Race, Education, and Occupation. Some caregivers started their caregiving for a family member, and later transitioned into professional caregiving. Serial Caregivers are marked with an ’*’ next to their ID.: table
@@ 3.2. Interview Procedure
We shared our screen with participants and asked them to rate eight caregiving-related concerns drawn from prior literature—1) disruptive patient behaviors (Swearer et al., 1988; Kunik et al., 2003; Desai et al., 2012), 2) insufficient support systems (Schulz and Martire, 2004; Jennings et al., 2015; Park et al., 2018), 3) doubt in self-efficacy (Tang and Chan, 2016; Gallagher et al., 2011; Crellin et al., 2014), 4) emotional wellness issue (Aminzadeh et al., 2007; Caputo, 2021; Gallego-Alberto et al., 2022), 5) relationship management (Ray, 2022; Vroman, 2019), 6) compassion fatigue (Day and Anderson, 2011; Day et al., 2014; Pérez et al., 2022), 7) no time for self-care (Waligora et al., 2019; Oliveira et al., 2019; Wang et al., 2019), and 8) burnout (Alves et al., 2019; Truzzi et al., 2012; Takai et al., 2009). We guided participants through each concern, providing literature-based definitions, and asked them to rate their level of concern on a scale from 1 (not at all concerning) to 5 (very concerning). Additionally, we asked the participants to explain their ratings and think aloud about specific experiences related to these concerns. This helped establish shared understanding and yielded deeper insights into participants’ mental wellbeing concerns which we qualitatively analyzed (Braun and Clarke, 2019). Table 3 summarizes participants’ responses to these prompts, where we find a high concern for most of the prompts. Finally, we asked the participants about the use of technology in caregiving as well as in managing mental wellbeing. Participants were encouraged to share their thoughts on the technologies they used, what features they found useful, and any suggestions or concerns they had for improving their experience. This helped us understand participants’ concerns and desires about these technologies.
@@ 3.3. Data Analysis
All co-authors participated in reviewing transcripts and engaged in an iterative open coding process, where codes were grouped, initial subthemes were identified, and these subthemes were refined into higher-level themes. The co-first authors led the coding effort, completing the majority of coding on raw interview transcripts, while the senior co-authors (who have extensive experience in qualitative research) coded two raw transcripts each during two two-hours hybrid (in-person and screen-sharing-based) co-working sessions. To ensure coherence, themes were reviewed, merged, split, or discarded based on relevance to the RQs, resulting in the final thematic structure. While we allowed flexibility in our coding and theme development, we employed thematic analysis (Braun and Clarke, 2019) as our primary methodological approach, guided by prior research on social support in online dementia communities (Johnson et al., 2022). A random subset of 10 transcripts was selected for an initial round of open coding. We developed 400 codes in the first iteration, which were then discussed by the entire team and distilled into 851 codes. The new set of codes were applied and iterated on the remaining transcripts. We discovered an additional 451 codes from the remaining transcripts. The codes were then grouped into 12 higher-level and 63 lower-level themes, which aligned with our three RQs.
@@ 4.1.1. Intersecting Sources of Caregivers’ Distress
Time Constraints and Limited Personal/Self-care Time. Participants expressed continual struggle with time management, as caregiving demands often left little room for self-care. Many felt overwhelmed, having to prioritize caregiving over their own wellbeing. P4 described caregiving as a highly time-consuming job, explaining that factors such as commuting, unexpected incidents, and the care recipient’s declining verbal abilities made it increasingly difficult to engage in other activities. P7 and P15 emphasized that caregiving often requires attention 24/7, leaving no time for breaks. Some, like P19, even changed jobs to better accommodate caregiving responsibilities:
“As a school Superintendent, that was really, really stressful. I was responsible for taking care of thousands of people and I couldn’t do that full time while also taking care of my mom full time. So I decided to resign and taking a job with the university where I can work from home.” —P19
@@ 4.1.2. Mental Health Impacts on Caregivers
The reality is that this is the only disease without a cure, and this disease is fatal, so we’re dealing a lot with people with anticipatory grief, which is what we talk about a lot. [..] And many times, people are not going to support groups until they’re like, literally in tears or at wit’s end. I see anger from men and frustration, and I see tears and physical demise from women. —P4
@@
“I feel increasing [caregiving] demands, my physical and emotional demand. This affected my mental health. I became more concerned about burnout and how to maintain my wellbeing.”—P21
@@
Overall, the caregiving-related concerns identified in prior research served as useful starting points for our conversations with participants (Table 3). These prompts helped guide participants to recall experiences related to emotional distress, burnout, and lack of support—all of which strongly resonated in their narratives. Many described feeling overwhelmed and neglecting their own wellbeing due to caregiving demands. Our findings also surfaced organically emerging themes, such as financial stress, loss of personal and social life, strained relationships, guilt around self-care, and limited family support—issues not always explicitly framed as mental health concerns in prior work. Notably, participants also shared self-reflective positives, including personal growth and emotional resilience, expanding the discourse beyond problem-focused narratives.
@@ 4.2. The Evolution of Mental Health of Caregivers
“When I first became a caregiver, I didn’t realize how much time and energy would be consumed by caregiving. I was in denial about how much this would impact on me emotionally.”—P21
@@
“I started to like experience, feelings of burnout and stress. I found out that I needed more emotional supports and not just from my colleagues, but also from therapy or peer groups.”—P17
@@
“Now, I have no one to open up to. I’m not comfortable sharing my problems to someone I know. I feel like they may judge me—or treat my mom differently when they visit our home.”—P3
@@
Late Stage: Emotional Exhaustion and Burnout. Over the significant progression of time and the decline of care recipient’s condition, caregivers adapt to their caregiving duties almost like a full-time job. However, by this point, they are also overwhelmed by providing 24/7 care, leaving little to no time for self-care. This period marks peak physical and emotional exhaustion, with burnout becoming a frequent concern. Participants mentioned severe sleep deprivation, as P11 described they would almost “sleep with one eye open.” P10 described the demands of managing every aspect of the care recipient’s life, including handling bills and estate matters, coordinating constant medical appointments, and providing hands-on care and support.
@@
Experience/Journey of Serial Caregivers. This three-stage process primarily reflects the typical experience of first-time caregivers. For serial caregivers—those who have cared for multiple family members with AD/ADRD, like P2, P16, P17, P21, and P22—the experience evolves across iterations. Technological advancements also shape these experiences; for instance, P2 noted that smart home technologies would have made a significant difference during their caregiving two decades ago.